At December 31,2019, the end of the annual reporting period, the accounts of Huron company showed the following:

1) Sales revenue for 2019, $2,950,000 of which one-quarter was on credit.

2) Allowance for doubtful accounts, balance January 1, 2019 is $22,700 credit.

3) Account receivable, December 31, 2019 (prior to any write-offs of uncollectible accounts during 2019, $383,400.

4) Uncollectible accounts to be written off, December 31,2019, $19,800. These accounts are all in the past due over 90 days category.

5) Aging schedule December 31,2019, showing the following break down of account receivable (prior to any write-offs of uncollectible accounts during 2019).

Status

Amount

Not past due

$210,800

Past due 1-60 days

$60,000

Past due over 60 days

$89,100

Past due over 90 days

$23,500

Required:

1) Give the 2019 entry to write off the uncollectible accounts.

2) Give the 2019 adjusting entry to record bad debt expense for each of the following independent assumptions concerning bad debt loss rates:

1) On credit sales. 1.8%

2) On total receivables at year-end (after write-off), 4.2%

3) On aging schedule;

Not past due

0.8%

Past due (1-60 days)

1.8%

Past due over 60 days

11%

Past due over 90 days

80%

3) Show the amount that would be reported on the 2019 statement of financial position relating to net account receivable for each assumption.
Answer #1

Entry for write-off of uncollectible accounts:

Bad debts account Debit 19800

Allowance for doubtful debts account Credit 19800

2)1)

Profit & Loss account Debit 13275

Allowance for doubtful debts account Credit 13275

2)2) Total receivables at year end = 383400.

Profit & Loss account Debit 16102.8

Allowance for doubtful debts account Credit 16102.8

2)3)

Calculation = 210800 * 0.8% + 60000 * 1.8% + 89100 * 11% + 23500 * 80% = 31367.4

Profit & Loss account Debit 31367.4

Allowance for doubtful debts account Credit 31367.4

3) The amounts that would be reported on the 2019 statement of financial position relating to net account receivable = 383400 - allowance for doubtful debts

For 1st assumption, reported figures showing net account receivable will be 383400 - 13275 = 370125

For 2nd assumption, reported figures showing net account receivable will be 383400 - 16102.8 = 367297.2

For 3rd assumption, reported figures showing net account receivable will be 383400 - 31367.4 = 352032.6